OpenStudy (anonymous):

need help solving sec(x/2)=cos(x/2) over the interval [0,2pi]. thx.

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos (x/2) } = \cos (x/2)\] \[\cos ^2 (x/2)=1\]

OpenStudy (anonymous):

Can you solve it further?